OpenStudy (anonymous):

simplify as much as possible: (81a^12)^1/4

OpenStudy (anonymous):

You can split 81 and a^12 into two separate quantities each to the 1/4th power. 81^(1/4) = 3 (a^12)^(1/4) = a^3 (81a^12)^1/4 = 3*a^3 You can check by raising 3*a^3 to the 4th power.